OpenStudy (anonymous):

use the second equation to make a substitution for x in the first equation. 3x + 2y = 7 x - y + 3 = 0 What is the resulting equation?

OpenStudy (anonymous):

What steps have you taken so far?

OpenStudy (anonymous):

+y-3. +y-3 X=y-3 3x+2y=7 3(y-3)+2y=7 3y-9+2y=7 5y-9=7

OpenStudy (here_to_help15):

3x + 2y = 7 x - y + 3 = 0 What is the resulting equation? x-y+3=0 +y-3. +y-3 X=y-3 3x+2y=7 3(y-3)+2y=7 3y-9+2y=7 5y-9=7

OpenStudy (anonymous):

thank you MR. here_to_help15